1. A company expects to sell 400 units of Product X in January and then expects sales to increase by 10% per month. If Product X sells for $10 each, the total sales for the first quarter of the year will be $

2. A manufacturing company expects to sell 12,000 units in August and 15,000 units in September. The company desires to have an ending inventory of 80% of the next month's sales. If inventory on August 1 is 8,000 units, then the company should produce units in August.

3. A manufacturer requires an ending inventory of 5,000 units. Their budgeted unit sales are 20,000 units and beginning finished goods inventor is 3,000 units. The units to be produced is

PLEASE EXPLAIN ME STEP BY STEP HOW DID YOU GOT THE RESULT
Answer #1

1) Total sales for first quarter

January February March Quarter
Sales in units 400 400*1.1 = 440 440*1.1 = 484 1324*10 = $13240

2) Production in August = 12000+(15000*80%)-8000 = 16000 Units

3) Production Unit = Sale unit+Desired ending inventory-beginning inventory = 20000+5000-3000 = 22000 Units